OpenStudy (anonymous):

what is the second derivative of y = 3/(4x^-3) a -3.6 b -1.84 c 2.4 d 4.5

OpenStudy (anonymous):

Well, what do you have to start out with? It's 3 * (4x^-3)^-1. Can you take this first derivative?

OpenStudy (phi):

re-write as \[y=\frac{3}{4}x^{3}\] Take the 2nd derivative, then plug in the x value they gave you (I assume x=1?)

OpenStudy (anonymous):

yes x =1

OpenStudy (anonymous):

kingkos, do you know how to find the second derivative?

OpenStudy (anonymous):

no not yet lesson for today of my prof

OpenStudy (anonymous):

oh okay. So did you see the way I worked the first derivative? I'll clean it up a little bit so it's more readable.

OpenStudy (phi):

do you know how to take the derivative? If so , do so

OpenStudy (anonymous):

yes thats the formula i write in my notebook but im confuse of your -4 how did u get that?

OpenStudy (anonymous):

\[[9\times(4x ^{-3}-3)^{-2}*(-12x ^{-4})\] I screwed up a bit on that last part in an earlier message, sorry. But I think this is correct.

OpenStudy (anonymous):

So you take the derivative of the outer part and then multiply by the derivative of the inner part.

OpenStudy (anonymous):

It just gets messier from here.

OpenStudy (anonymous):

crap, I pasted in the equation, but the whole thing isn't showing up.

OpenStudy (anonymous):

nevermind! My brain isn't working AT ALL today. Yeah, it's 4.5

OpenStudy (phi):

I believe y = 3/(4x^-3) means \[ y=\frac{3}{4x^{-3}} \] which is equivalent to \[ y=\frac{3}{4} x^{3} \] Use the power rule \[ \frac{dy}{dx}= \frac{9}{4} x^2 \] the 2nd derivative is \[ \frac{d^{2}y}{dx^2}= \frac{18}{4} x \] Evaluate at x=1 to find 4.5

OpenStudy (anonymous):

sorry phi and kingkos, I'm a total retard today!

OpenStudy (phi):

Everyone has a bad hair day every now and again.

OpenStudy (anonymous):

This is what happens when I visit this site early in the morning, LOL

OpenStudy (anonymous):

things phi dont worry i after i finish my homework i will study hard

OpenStudy (anonymous):

its ok brinethery same with me